OpenStudy (anonymous):

Find real number a and b such that : (3+2i)a+(4+5i)b=18+19i.

OpenStudy (anonymous):

\((3+2i)a+(4+5i)b=(3a+4b)+(2a+5b)i=18+19i\) \( \iff 3a+4b=18 \text{ and } 2a+5b=19.\)

OpenStudy (anonymous):

Solve the two equations to get a and b.

OpenStudy (anonymous):

a=2 , b=3.